given f(x)=20x+7 and g(x) =2x^2-4x+13, find (f-g)(x) then evaluate the difference when x=10.
Answer by stanbon(75887)

given f(x)=20x+7 and g(x) =2x^2-4x+13, find (f-g)(x) then evaluate the difference when x=10.
----
(f-g)(x) = f(x)-g(x)
---
= 20x+7 - (2x^2-4x+13)
----
= -2x^2 +24x -6
-----
(f-g)(10) = -2(10^2)+24(10)-6
= -200 + 240 -6
---
= 34
=====
Cheers,
Stan H.